# EyeEcho: Continuous and Low-power Facial Expression Tracking on Glasses

### 13 Feb 2024 v1

@@ Abstract.
ccs: Human-centered computing Ubiquitous and mobile devicesccs: Hardware Power and energyjournalyear: 2024copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642613isbn: 979-8-4007-0330-0/24/05
@@ 3.2. Principle of Operation
Prior work (Li et al., 2022; Wu et al., 2021; Zhang et al., 2023c; Chen et al., 2020) have demonstrated that partial skin and muscle deformations behind and inside ears are highly informative to reconstruct full facial expressions when they are captured by different kinds of sensors. Xie et al. (Xie et al., 2021) proved that the skin and muscle deformations around the eyes and the cheeks contain information that can be extracted to recognize upper facial gestures. The sensing hypothesis of EyeEcho is that these deformations around eyes and cheeks are highly informative about detailed facial movements on both lower and upper face including eyes, eyebrows, cheeks, and mouth. Considering that people’s facial skin and muscles are interconnected, moving any part of the face would inevitably stretch the muscles and skin on the entire face. EyeEcho applies this skin-deformation-based sensing hypothesis on glasses to develop research questions: is it possible to infer full facial expressions by only observing skin deformations around glasses (e.g., cheeks)? If so, what is the appropriate hardware set up including the number, orientation and position of the sensors? How well can it track facial movements on different areas of the face (e.g., blinking) under different real-world scenarios? To explore these research questions, we developed EyeEcho using active acoustic sensing to capture the skin deformations on glasses, which we will detail later. We chose acoustic sensing because its sensors are small, lightweight, low-power and have been successfully applied to various tasks on tracking human activities, including health-related activities detection (Nandakumar et al., 2015; Wang et al., 2018), novel interaction methods (Aumi et al., 2013; Wang et al., 2016; Yun et al., 2017; Xu et al., 2020), silent speech recognition (Zhang et al., 2020, 2023b, 2023a; Sun et al., 2023), authentication (Lu et al., 2019; Gao et al., 2019; Wang et al., 2022; Huh et al., 2023; Iwakiri and Murao, 2023), discrete facial expression recognition (Xie et al., 2021), finger tracking (Sun et al., 2018; Nandakumar et al., 2016), hand gesture recognition (Zhang et al., 2018b), body pose estimation (Mahmud et al., 2023), and motion tracking (Zhang et al., 2018a; Lian et al., 2021).
@@ 5. Evaluation of EyeEcho in an In-lab Study
Figure 5. The Degree of Deformation of Different Facial Parts when Performing Different Facial Expressions.
@@ 5.1.2. Selection of Facial Expressions for Continuous Tracking
To emphasize this point, we initially defined the "degree of deformation of one facial part" as the average of the blendshape parameters associated with that specific part out of the 52 blendshape parameters provided by the ARKit API, as introduced in Sec. 4.3.1. Subsequently, we plotted the degree of deformation of different facial parts while participants performed these nine facial expressions, as depicted in Fig. 5. As previously discussed in Sec. 4.3.1, theoretically, the maximum possible degree of deformation for one facial part is 1000. The figure illustrates that nearly all facial parts, with the exception of the tongue (related to only one blendshape parameter), exhibit some degree of movement when performing either upper or lower facial expressions. This observation underscores the interconnected nature of facial muscles. Moreover, the movements of the mouth, cheeks, and nose are even more pronounced when upper facial expressions, such as "Close Eyes," are performed. This substantiates our claim that these nine selected facial expressions effectively evaluate our system’s performance in tracking multiple facial parts simultaneously. It’s worth noting that these expressions have also been utilized in prior studies, such as (Li et al., 2022). During the study, participants were presented with these facial expressions multiple times in an instructional video to mimic, and the order of presentation was randomized in each session to mitigate the influence of expression sequence.
@@ 5.4. Performance on Tracking Facial Expressions with Varying Degrees of Deformation
Since our EyeEcho system tracks facial expressions continuously rather than classifying them, we capture the entire process of a facial expression transitioning from a neutral face to its most extreme state. In the previous subsection, we validated the overall performance of our EyeEcho system. However, it is also important to assess its performance in tracking different degrees of deformation for these facial expressions. Based on our analysis in Sec. 5.1.2, we calculated the degree of deformation for each frame and categorized all frames in the user study into four groups based on their degrees of deformation. We present the evaluation results in terms of average MAE for each category in Tab. 3, along with the percentage that each category represents among all the frames. As shown in the table, the majority of frames in our user study have deformation levels below 150. Our system demonstrates satisfactory performance on tracking facial expressions for these frames, as Subsection 5.2 establishes that the prediction and ground truth are visually highly similar when the MAE is below 40. For frames with deformation levels exceeding 150, the performance of our system is slightly lower, but it still maintains an acceptable level of accuracy, as depicted in Fig. 9. This analysis demonstrates that our system performs well across varying degrees of deformation for different facial expressions.
@@ 5.5. Determining Minimum Training Data Requirement
In the previous experiments, we used the data from 10 sessions (20 minutes of data) to train the model. However, 20 mins of training may not be always preferable for users in real-world deployments. Therefore, we further explore how much training data is needed before the system reaches an acceptable accuracy. We chose two sessions of data as the testing sessions and employed 2, 4, 6, 8 and 10 sessions of data respectively to train the model for each scenario. We showed the evaluation results in Fig. 10 (a). As we can see in the figure, the average performance of the system improved with more training data. However, even with just 2 training sessions (4 minutes of training data), the system achieved a MAE of 29.7 and 34.3 for the two scenarios. According to the analysis in Sec. 5.2 and Sec. 5.3.2, this result is already good enough to provide an acceptable tracking performance on facial expressions that are highly similar to the ground truth visually. In essence, two sessions of training data are likely enough to provide satisfactory tracking performance in real-world deployments for the majority of users. More training sessions could be collected for users who do not have good enough performance.
(a) MAE vs NT
(b) MAE vs Models
Figure 10. Impact on MAE with Different Number of Training Sessions (NT) and Models.
@@ 5.6. User-Adaptive Model
The results of this user-independent model are presented in Figure 10 (b). Notably, in comparison to the results of the user-dependent model, the results were worse, with a MAE of 49.0 and 53.2 for the sitting and walking scenarios, respectively. However, this outcome was anticipated because our system relies on the reflection of acoustic signals on the face and head, which varies significantly among participants. Additionally, differences in how participants wore the device and executed facial expressions also contributed to this variance.
@@ 6.1. Impact of Operating Frequency Range
6.2. Impact of Environmental Noises
@@ 7.1. Study Design
(a) Living Room
(b) Bedroom
(c) Kitchen
Figure 16. Apartment used for the Semi-in-the-wild Study.
@@ 7.3. Study Procedure
(a) MAE vs. Participants (without fine-tuning)
(b) MAE vs. Experiment Rooms (without and with fine-tuning)
Figure 17. Performance of EyeEcho for the Semi-in-the-wild Study.

### 23 Feb 2024 v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the CHI Conference on Human Factors in Computing Systems; May 11–16, 2024; Honolulu, HI, USAbooktitle: Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24), May 11–16, 2024, Honolulu, HI, USAdoi: 10.1145/3613904.3642613isbn: 979-8-4007-0330-0/24/05ccs: Human-centered computing Ubiquitous and mobile devicesccs: Hardware Power and energy
@@ 1. Introduction
This figure shows the nine facial expressions used in our study performed by one person wearing the glasses form factor embedded with our EyeEcho system, which are Open Mouth, |o| Mouth, Open Eyes, Close Eyes, Close Left Eye, Close Right Eye, Smile Face, Sneer Left, and Sneer Right. Under the nine facial expressions, the figure shows an avatar with these nine facial expressions which are predicted and generated by our EyeEcho system. Then under these are nine figures of different echo profiles corresponding to these nine facial expressions for both left and right channels of microphones. The differential echo profiles display different patterns for different facial expressions and are used for model prediction. On the right of the figure, a picture of our EyeEcho device and a person wearing it are shown.
@@ 3.2. Principle of Operation
Prior work (Li et al., 2022; Wu et al., 2021; Zhang et al., 2023c; Chen et al., 2020) have demonstrated that partial skin and muscle deformations behind and inside ears are highly informative to reconstruct full facial expressions when they are captured by different kinds of sensors. Xie et al. (Xie et al., 2021) proved that the skin and muscle deformations around the eyes and the cheeks contain information that can be extracted to recognize upper facial gestures. The sensing hypothesis of EyeEcho is that these deformations around eyes and cheeks are highly informative about detailed facial movements on both lower and upper face including eyes, eyebrows, cheeks, and mouth. Considering that people’s facial skin and muscles are interconnected, moving any part of the face would inevitably stretch the muscles and skin on the entire face. EyeEcho applies this skin-deformation-based sensing hypothesis on glasses to develop research questions: is it possible to infer full facial expressions by only observing skin deformations around glasses (e.g., cheeks)? If so, what is the appropriate hardware set up including the number, orientation and position of the sensors? How well can it track facial movements on different areas of the face (e.g., blinking) under different real-world scenarios? To explore these research questions, we developed EyeEcho using active acoustic sensing to capture the skin deformations on glasses, which we will detail later. We chose acoustic sensing because its sensors are small, lightweight, low-power and have been successfully applied to various tasks on tracking human activities, including health-related activities detection (Nandakumar et al., 2015; Wang et al., 2018), novel interaction methods (Aumi et al., 2013; Wang et al., 2016; Yun et al., 2017; Xu et al., 2020), silent speech recognition (Zhang et al., 2020, 2023b, 2023a; Sun et al., 2023), authentication (Lu et al., 2019; Gao et al., 2019; Wang et al., 2022; Huh et al., 2023; Iwakiri and Murao, 2023), discrete facial expression recognition (Xie et al., 2021), gaze tracking (Li et al., 2024), finger tracking (Sun et al., 2018; Nandakumar et al., 2016), hand gesture recognition (Zhang et al., 2018b; Lee et al., 2024), body pose estimation (Mahmud et al., 2023), and motion tracking (Zhang et al., 2018a; Lian et al., 2021).
@@ 4.1. Hardware Prototype Design
Figure 2. Design of Hardware and Glasses Form Factor.
@@
This figure features four subfigures. Subfigure (a) shows the speakers, microphones and the BLE module on our customized PCBs, which are used in our system. The BLE module is 20 mm in length. Subfigure (b) displays a pair of glasses and it is marked with three positions, i.e., Position 1 on the legs of the glasses, Position 2 on the bottom of the glass frame and Position 3 is under the bridge of the glass frame. Subfigure (c) shows the complete prototype of the glasses embedded with our EyeEcho system. A pair of speaker and microphone is placed on each leg of the glasses. A 3D printed case is placed on the left leg of the glasses housing the BLE module and the battery. Subfigure (d) shows a person wearing the glasses prototype, demonstrating the position of the speakers, microphones, and the case relative to the user’s face.
@@ 4.1.3. Form Factor Design
This figure displays the overview of the EyeEcho system. Subfigure (a) shows a person wearing the glasses prototype. Subfigure (b) shows the PCBs of speakers, microphones and BLE module. Subfigure (c) shows the transmitted FMCW signal in time-frequency domain. This signal sweeps from 16-20 kHz. Subfigure (d) shows the signal reflected from the facial skin. Subfigure (e) shows the received FMCW signal in time-frequency domain. The signal sweeps from 16-20 kHz but has some noises in it. Subfigure (f) shows a Butterworth Band-pass filter from 16-20 kHz in the frequency domain. Subfigure (g) shows the echo profile of one facial expression. It displays the unique patterns associated with the facial expression. Subfigure (h) shows the differential echo profile which more clearly demonstrates the pattern of the facial expression. Subfigure (i) shows a deep learning module with the ResNet-34 architecture. Subfigure (j) shows the avatar with the estimated facial expressions corresponding to the facial expressions that the person performs.
Figure 3. Overview of EyeEcho System.
@@ 4.2.2. Echo Profile Calculation
This figure shows the differential echo profiles of three facial movements in both left channel and right channel. For Sneer Left and Sneer Right, there is a clear pattern in the corresponding channel in the differential echo profiles. There is a weaker pattern in the other channel. For subtle facial movement like blinking, the pattern is weaker and shorter which aligns with nature of this movement.
@@ 5. Evaluation of EyeEcho in an In-lab Study
Figure 5. The Degree of Deformation of Different Facial Parts when Performing Different Facial Expressions to the Most Extreme State.
This figure shows the degree of deformation of different facial parts, including eye, eyebrow, jaw, mouth, cheek, nose and tongue, when the nine facial expressions and blinking are performed. When any of the facial expressions is performed, all facial parts deform to a certain extent for both lower and upper facial parts.
Figure 5. The Degree of Deformation of Different Facial Parts when Performing Different Facial Expressions to the Most Extreme State.
@@ 5.1.2. Selection of Facial Expressions for Continuous Tracking
To emphasize this point, we initially defined the "degree of deformation of one facial part" as the average of the blendshape parameters associated with that specific part out of the 52 blendshape parameters provided by the ARKit API, as introduced in Sec. 4.3.1. Subsequently, we plotted the degree of deformation of different facial parts while participants performed these nine facial expressions to the most extreme state, as depicted in Fig. 5. As previously discussed in Sec. 4.3.1, theoretically, the maximum possible degree of deformation for one facial part is 1000. The figure illustrates that nearly all facial parts, with the exception of the tongue (related to only one blendshape parameter), exhibit some degree of movement when performing either upper or lower facial expressions. This observation underscores the interconnected nature of facial muscles. Moreover, the movements of the mouth, cheeks, and nose are even more pronounced when upper facial expressions, such as "Close Eyes," are performed. This substantiates our claim that these nine selected facial expressions effectively evaluate our system’s performance in tracking multiple facial parts simultaneously. It’s worth noting that these expressions have also been utilized in prior studies, such as (Li et al., 2022). During the study, participants were presented with these facial expressions multiple times in an instructional video to mimic, and the order of presentation was randomized in each session to mitigate the influence of expression sequence. In Fig. 5, we also included a subtle facial movement, blinks, because they happened spontaneously to the participants in the study.
@@
Figure 6. Visualization of Varying Degrees of Deformation.
This figure shows the visualization of four facial expressions, Open Mouth, Smile Face, Open Eyes, and Close Eyes, with different degrees of deformation. For Open Mouth, the face reaches the most extreme state with 100 Deformation while for the rest three facial expressions, the face reaches the most extreme state with 150 Deformation.
@@ 5.1.3. Study Procedure
This figure visualizes the prediction of three facial expressions with different values of MAE compared with ground truth, which are Open Mouth, Open Eyes and Blink. For large facial expressions like Open Mouth and Open Eyes, the prediction is visually similar to ground truth when MAE is 40. When MAE is around 20, the prediction is basically indistinguishable from the ground truth. For subtle facial expressions like blinking, when the MAE is around 20, the prediction is also visually highly similar to the ground truth.
@@ 5.3. User-Dependent Model
This figure shows the Cumulative Distribution Function (CDF) of MAE for all participants. Subfigure (a) for sitting scenario and Subfigure (b) for walking scenario. Generally, LMAE is smaller than UMAE.
@@ 5.3.1. Numerical Results
This figure shows the visualization of two facial expressions, Smile Face and Close Eyes. For each facial expression, we selected three frames and displayed both the ground truth expression and the predicted facial expression, along with the corresponding MAE/LMAE/UMAE. These frames were randomly selected from different participants. The prediction and the ground truth are visually highly similar to each other when LMAE is under 40 and UMAE is under 60.
@@ 5.4. Performance on Tracking Facial Expressions with Varying Degrees of Deformation
Since our EyeEcho system tracks facial expressions continuously rather than classifying them, we capture the entire process of a facial expression transitioning from a neutral face to its most extreme state. In the previous subsection, we validated the overall performance of our EyeEcho system. However, it is also important to assess its performance in tracking different degrees of deformation for these facial expressions. Fig. 5 depicts the degree of deformation of different facial parts in frames where the facial expressions are performed to the most extreme state but they usually only account for a small portion of all frames in the entire process of making facial expressions because this is a continuous process from the neutral face to the most extreme state of facial expressions and then back to the neutral face. To quantify this proportion, based on our analysis in Sec. 5.1.2, we calculated the degree of deformation for each frame and categorized all frames in the user study into four groups based on their degrees of deformation. We present the evaluation results in terms of average MAE for each category in Tab. 3, along with the percentage that each category represents among all the frames.
As shown in the table, the majority of frames in our user study have deformation levels below 150, in alignment with what has been shown in Fig. 6. The frames plotted in Fig. 5 mostly fall into the last category in the table with the degree of deformation larger than 150, taking up less than 10% of all the frames. Note that the maximum degree of deformation that the subtle facial movement blinking can reach is usually below 100, as shown in Fig. 5. Our system demonstrates satisfactory performance on tracking facial expressions for the frames with a degree of deformation smaller than 150, as Subsec. 5.2 and Fig. 7 establish that the prediction and ground truth are visually highly similar when the MAE is below 40. For frames with deformation levels exceeding 150, the performance of our system is slightly lower, but it still maintains an acceptable level of accuracy, as depicted in Fig. 7 and Fig. 9. For blinking which has more subtle degree of deformation (smaller than 100), our system also achieves promising tracking performance considering that the MAE is around 20, as compared to the visualization in Fig. 7. This analysis demonstrates that our system performs well across varying degrees of deformation for different facial expressions.
@@ 5.5. Determining Minimum Training Data Requirement
In the previous experiments, we used the data from 10 sessions (20 minutes of data) to train the model. However, 20 mins of training may not be always preferable for users in real-world deployments. Therefore, we further explore how much training data is needed before the system reaches an acceptable accuracy. We chose two sessions of data as the testing sessions and employed 2, 4, 6, 8 and 10 sessions of data respectively to train the model for each scenario. We showed the evaluation results in Fig. 10. As we can see in the figure, the average performance of the system improved with more training data. However, even with just 2 training sessions (4 minutes of training data), the system achieved a MAE of 29.7 and 34.3 for the two scenarios. According to the analysis in Sec. 5.2 and Sec. 5.3.2, this result is already good enough to provide an acceptable tracking performance on facial expressions that are highly similar to the ground truth visually. In essence, two sessions of training data are likely enough to provide satisfactory tracking performance in real-world deployments for the majority of users. More training sessions could be collected for users who do not have good enough performance.
Figure 10. Impact on MAE with Different Number of Training Sessions (NT).
This figure shows the MAE with different number of training sessions. It shows that for both sitting and walking, more training sessions lead to better performance.
Figure 10. Impact on MAE with Different Number of Training Sessions (NT).
@@ 5.6. User-Adaptive Model
The results of this user-independent model are presented in Fig. 11. Notably, in comparison to the results of the user-dependent model, the results were worse, with a MAE of 49.0 and 53.2 for the sitting and walking scenarios, respectively. However, this outcome was anticipated because our system relies on the reflection of acoustic signals on the face and head, which varies significantly among participants. Additionally, differences in how participants wore the device and executed facial expressions also contributed to this variance.
Figure 11. Impact on MAE with Different Models.
This figure shows the MAE with different ML models. It shows that user-independent model leads to worst performance while user-adaptive model with 2 training sessions yields better performance than user-dependent model with the same amount of training data.
Figure 11. Impact on MAE with Different Models.
@@ 5.7. Transfer Learning using Data from Sitting to Walking Scenario
This figure demonstrates the average MAE for all 10 participants in the study with the operating frequency range set as 20-24 kHz. They are 27.6, 15.3, 31.4, 18.7, 13.2, 22.9, 20.3, 24.1, 20.5, 42.4 for P1-P10. The average is 23.6.
@@ 6.1. Impact of Operating Frequency Range
Figure 13. Three Different Noisy Environments.
This figure shows three noisy environments. (a) is a quiet meeting room with chairs and tables, (b) is a cafe area with many people working and talking and (c) is a crossroad on the street with vehicles and pedestrians passing by.
@@ 6.2. Impact of Environmental Noises
This figure shows the signal in our system together with the cafe noise in the frequency domain. Besides, it also shows three noises, which are cafe noise, music noise and street noise, in the frequency domain. The noises are mainly in the audible frequency range (under 5 kHz) and the strength of them are much smaller compared to the signal.
@@
This figure displays the signal with no noise, music noise, cafe noise and street noise in the frequency domain from P5’s data. The first three signals look very similar while the shape of the signal with street noise is very different.
The figure shows the signal change after the device was moved from indoor to outdoor. 1 minute after the device was moved outdoor, the signal shape changed. 2 minutes and 3 minutes after, the signal shape continued to change and became stable.
@@ 6.3. Usability of EyeEcho
Figure 17. Apartment used for the Semi-in-the-wild Study.
(a) Living Room
(b) Bedroom
(c) Kitchen
This figure shows the furniture layout of three rooms, living room, bedroom and kitchen in a one-bedroom apartment.
Figure 17. Apartment used for the Semi-in-the-wild Study.
@@ 7.2. Study Environment and Activities
Figure 18. Performance of EyeEcho for the Semi-in-the-wild Study.
(a) MAE vs. Participants (without fine-tuning)
(b) MAE vs. Experiment Rooms (without and with fine-tuning)
This figure shows performance of our system for the semi-in-the-wild study. Subfigure (a) shows MAE without fine-tuning for all 10 participants and average in the living room, bedroom and kitchen. Living room has the best performance and the kitchen has the worst performance. Generally, Day 1 has better performance than Day 2. Subfigure (b) shows MAE with and without fine-tuning. Fine-tuning improves performance for all three rooms and both days, especially for kitchen.
Figure 18. Performance of EyeEcho for the Semi-in-the-wild Study.
@@ 7.4.3. Results
This figure shows the visualization of results in the semi-in-the-wild study. We displayed the performance for two activities, watching videos and talking, and selected three frames for each activity. The frames were randomly selected from all participants. The frames are marked with corresponding MAE/LMAE/UMAE. This figure shows that the prediction and the ground truth are visually similar to each other when LMAE is under 40 and UMAE is under 60.
@@ 8.1. Power Consumption Analysis
This figure shows two speakers side by side. They are soldered on the same type of PCB and very similar to each other. OWR-05049T-38D is slightly smaller than SR6438NWS-000.
@@ 8.2. Real-time Deployment on Smartphones
This figure shows the real-time pipeline implemented on an Android phone. In the figure, a person is wearing the glasses prototype and doing the Open Mouth facial expression. The model deployed on the phone is making predictions in real-time. On the screen of a laptop, the personalized avatar is displaying the same facial expression that the person is performing.